# Constraints on SMEFT operators from $Z\to\mu\mu bb$ decay

### v1

@@ Abstract
The Standard Model Effective Field Theory (SMEFT) provides a systematic framework to probe indirect effects of heavy new physics via precision measurements. While SMEFT constraints have been extensively studied using purely leptonic $Z$ decays and inclusive $Z$ production, mixed leptonic–hadronic modes remain largely unexplored. In this work, we analyze $Z\to\mu\mu bb$ decays within the SMEFT framework, deriving constraints on dimension-six operators that affect four-fermion interactions between leptons and bottom quarks, as well as ZZ–fermion couplings. Signal and background events are simulated with state-of-the-art Monte Carlo tools, including detector effects such as $b$-tagging, and limits on the relevant Wilson coefficients are extracted using kinematic distributions and a profile likelihood approach. Our results provide complementary constraints to existing SMEFT studies and yield the first process-specific limits on flavor-resolved four-fermion operators involving muons and bottom quarks from $Z$ decays.
@@ I Introduction
The Standard Model Effective Field Theory (SMEFT) offers a model-independent framework to parameterize the effects of heavy new physics through higher-dimensional operators constructed from SM fields [4, 5, 6]. In recent years, SMEFT interpretations of $Z$-boson observables have attracted considerable attention. Purely leptonic decay modes, such as $Z\to 4\ell$, provide stringent constraints on electroweak and four-lepton operators, while complementary information has been obtained from SMEFT analyses of inclusive $Z$ production and ZZ+jets, probing quark electroweak couplings and dipole interactions [7, 8, 9, 10].
Despite this progress, $Z$-boson decays into mixed leptonic–hadronic final states remain relatively unexplored in the SMEFT context. Such channels are sensitive to a distinct set of operators, in particular four-fermion contact interactions involving both leptons and quarks, as well as flavor-dependent modifications of ZZ–fermion couplings [11]. Among these, the decay $Z\to\mu\mu bb$ is of special interest. The presence of bottom quarks enhances sensitivity to operators coupling to the third quark generation, while the dimuon final state allows for precise reconstruction of the leptonic kinematics. At the same time, the analysis poses experimental challenges due to sizable QCD backgrounds and the reliance on heavy-flavor tagging [12, 13].
Figure 1: Representative Feynman diagram contributing to the process $Z\to\mu\mu bb$ in the Standard Model, where the four-body final state arises from final-state radiation of an off-shell $Z^{\ast}/\gamma^{\ast}$.
@@
In this work, we perform a dedicated SMEFT analysis of the decay $Z\to\mu\mu bb$ and derive constraints on dimension-six operators contributing to this process. We focus on operators that induce four-fermion interactions of the $\ell\ell bb$ type and on operators that modify the $Z$ couplings to leptons and bottom quarks. Signal and background processes are simulated using state-of-the-art Monte Carlo tools, and detector-level effects are included [14, 15, 16]. Limits on the WCs are obtained from differential distributions using a profile likelihood approach [17].
This study provides a complementary probe of new physics effects in the $Z$-boson sector, extending existing SMEFT constraints beyond purely leptonic final states. It demonstrates the potential of mixed leptonic–hadronic $Z$ decays to access operator directions that are weakly constrained by current analyses and motivates their inclusion in future global SMEFT fits [18].
@@ II Effective field theory description
where $\mathcal{O}_{i}$ are dimension-six operators and $C_{i}$ the corresponding WCs [10, 11]. In this work, we truncate the expansion at dimension six contributions [19, 20].
@@
The decay $Z\to\mu\mu bb$ receives contributions from several classes of dimension-six operators. First, four-fermion operators involving leptons and bottom quarks generate direct contact interactions contributing to the four-body final state at tree level [7, 8]. Schematically, such operators can be written as $(\bar{\ell}\Gamma\ell)(\bar{q}\Gamma q)$ and induce amplitudes that interfere with the SM contributions mediated by off-shell electroweak bosons.
@@
Second, operators that modify the couplings of the $Z$ boson to fermions lead to indirect contributions to the decay process. After electroweak symmetry breaking, these operators generate shifts in the effective ZZ–lepton and ZZ–quark couplings, which affect both the production and decay of the $Z$ boson [21, 18]. These effects propagate into the $Z\to\mu\mu bb$ final state through the modified electroweak vertices.
@@
In the numerical analysis, we select a set of six dimension-six operators that contribute to this decay channel and study their impact individually. The operators are implemented using the SMEFTsim_general_MwScheme_UFO model [6, 22], which is compatible with MadGraph5_aMC@NLO [14] and allows event generation and reweighting at leading order in a consistent electroweak input scheme.
In this analysis, we retain both the interference and quadratic terms in the WCs when evaluating SMEFT effects [7]. The dependence of physical observables on a given WC is therefore parameterized by a quadratic function, which can be efficiently determined using the reweighting technique [14]. This approach allows us to explore the parameter space of the WCs without generating independent event samples for each point.
The validity of the SMEFT expansion relies on the assumption that the typical energy scale of the process remains below the cutoff scale $\Lambda$ [19, 10]. In practice, the absence of events with extremely large $m_{\mu\mu bb}$ or jet transverse momenta indicates that the analysis is dominated by energy scales safely below the assumed new physics cutoff.
@@ III Event generation and simulation
All processes are generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8 [15]. Detector-level effects are simulated using Delphes [16] with an CMS-like detector configuration. Unless stated otherwise, proton–proton collisions at a center-of-mass energy of $\sqrt{s}=13~\mathrm{TeV}$ are assumed throughout this analysis.
The signal process considered in this work is the four-body decay of the Z boson into a dimuon pair and a bottom–antibottom quark pair: $Z\to\mu\mu bb$. At leading order, the four-body final state arises predominantly from final-state radiation or off-shell electroweak boson exchange [7]. Representative Feynman diagrams contributing to the signal process are shown in Fig. 1.
Signal events are generated at parton level using MadGraph5_aMC@NLO. Basic generator-level cuts are applied to ensure numerical stability and reflect the kinematic acceptance of the detector: $p_{T}^{\ell}>10~\mathrm{GeV}$, $|\eta^{\ell}|<2.5$ for charged leptons, and $p_{T}^{q}>10~\mathrm{GeV}$, $|\eta^{q}|<2.5$ for quarks. The resulting leading-order cross section for the signal process is $\sigma(Z\to\mu^{+}\mu^{-}b\bar{b})=0.750~\mathrm{fb}$, where the quoted uncertainty corresponds to the Monte Carlo integration error.
The reconstructed kinematic distributions shown in Fig. 2 provide an important validation of the signal modeling after detector effects are taken into account. In Fig. 2(a), the invariant mass distribution of the $b\bar{b}$ system is considerably broader than that of dimuon system, as expected from jet energy resolution effects and QCD radiation. Nevertheless, the combined four-body invariant mass $m_{\mu\mu bb}$ displays a clear enhancement around $m_{Z}$, indicating that the signal topology can be reliably reconstructed at detector level despite the presence of hadronic final states.
@@
Figure 2: Detector-level kinematic distributions for the process $Z\to\mu\mu bb$ obtained from Delphes simulation.
@@
Subleading backgrounds from electroweak processes involving photons are also considered, including $Z\gamma$, $\gamma\gamma$, and mixed $\gamma Z$ production, where the photon converts into a fermion pair.
@@
Table 1: Leading-order cross sections for the signal and background processes considered in this analysis. The quoted uncertainties correspond to Monte Carlo integration errors. The merged ZZ+jets sample, combining matrix-element calculations with up to two additional jets, is used in the final analysis, while the individual samples are shown for reference.
@@ IV Event Reconstruction and Selection Strategy
Bottom-quark jets are identified using the $b$-tagging information stored in the Delphes output [16, 29]. A jet is considered a $b$-jet if its $b$-tag discriminator exceeds a threshold of $0.5$, corresponding to a typical medium working point used in CMS analyses. Jets are identified as $b$-jets using the $b$-tagging discriminator provided by Delphes, adopting a medium working point.
@@
To suppress backgrounds containing genuine missing transverse energy, such as top-quark pair production, events are required to have $\mathrm{MET}<30~\mathrm{GeV}$. This exploits the absence of neutrinos in the signal process $Z\to\mu\mu bb$.
@@
The selection criteria are designed to preserve the characteristic kinematics of the $Z\to\mu\mu bb$ decay while suppressing backgrounds with genuine missing energy or misidentified heavy-flavor jets. For events passing all selection requirements, the invariant mass of the $\mu\mu bb$ system is reconstructed using the four-momenta of the two leading isolated muons and the two selected $b$-jets. The reconstructed invariant mass $m_{\mu\mu bb}$, serves as the primary observable for signal extraction and the subsequent SMEFT analysis [7].
The signal region is defined as $80~\mathrm{GeV}<m_{\mu\mu bb}<115~\mathrm{GeV}$, centered around the $Z$ boson mass. This window retains the majority of the signal while efficiently suppressing non-resonant backgrounds. This window is chosen to maximize sensitivity to SMEFT-induced rate deviations while minimizing contamination from non-resonant backgrounds.
@@
FIG. 3(b) focuses on the signal region defined by $80<m_{\mu\mu bb}<115$ GeV, which is centered around the $Z$ boson mass. In this window, the signal contribution becomes clearly distinguishable, while the $t\bar{t}$ background is strongly suppressed by the missing transverse energy requirement. The remaining backgrounds, including $ZA$ and $ZH$ production, exhibit similar kinematic features but are subleading in rate. This demonstrates that the chosen invariant mass window provides an optimal balance between signal efficiency and background rejection, forming a robust basis for the subsequent SMEFT analysis. The clear localization of the signal contribution within the $Z$-mass window demonstrates that the invariant mass $m_{\mu\mu bb}$ provides a powerful discriminator for isolating SMEFT effects in this channel. This observable is therefore chosen as the primary input for the subsequent reweighting and limit-setting procedure.
V SMEFT Reweighting and Constraining
The impact of dimension-six operators on the process $Z\to\mu\mu bb$ is evaluated within the SMEFT framework using an event reweighting technique [31, 32]. In this approach, the effect of each WC on the cross section and kinematic distributions is obtained by reweighting SM Monte Carlo events. The total event weight includes the Standard Model contribution as well as linear and quadratic corrections from dimension-six operators, allowing an efficient exploration of the WCs parameter space without regenerating new event samples.
SMEFT effects are incorporated using the SMEFTsim_general_MwScheme_UFO model from the SMEFTsim 3.0 framework [6, 22], which provides a complete and gauge-invariant implementation of dimension-six operators in the $M_{W}$ input scheme. The $M_{W}$ scheme is adopted to consistently account for electroweak input-parameter shifts induced by dimension-six operators. For each generated event, a set of precomputed weights corresponding to different values of $\theta$ is stored and propagated through parton showering, hadronization, and detector simulation, ensuring a consistent treatment of SM and SMEFT contributions [15, 16].
@@ V SMEFT Reweighting and Constraining
where $c$ denotes the Standard Model (SM) expectation, $b$ corresponds to the interference between the SM and dimension-six amplitudes, and $a$ encodes the pure quadratic SMEFT contribution [7]. The coefficients $a$, $b$, and $c$ are determined by fitting the event yields obtained from reweighted samples generated at several fixed values of $\theta$.
@@
which, under the assumptions of Wilks’ theorem, follows a $\chi^{2}$ distribution with one degree of freedom [33]. The 95% confidence level (C.L.) interval on $\theta$ is obtained by requiring $q(\theta)=3.84$.
The resulting constraint is quoted both in terms of the WCs, $\theta_{95}=C_{i}/\Lambda^{2}$, and in terms of the relative deviation of the expected event yield with respect to the SM prediction,
$R_{95}\equiv\frac{N(\theta_{95})}{N_{\mathrm{SM}}}.$ (5)
This procedure is applied independently to each dimension-six operator, with all other WCs set to zero, providing a measure of the sensitivity of the $Z\to\mu\mu bb$ channel to physics beyond the Standard Model within the SMEFT framework. In this analysis, a flavor-specific setup is adopted, in which the WCs are defined for individual fermion generations. In particular, operators involving second-generation leptons ($\ell_{2}$) and third-generation quarks ($q_{3}$) are considered explicitly. This choice allows a direct probe of flavor-resolved SMEFT effects in the $Z\to\mu\mu bb$ final state, beyond the flavor-universal assumptions commonly employed in global SMEFT fits.
@@
The observed quadratic behavior confirms the validity of the reweighting procedure and illustrates the relative importance of linear and quadratic terms for different operator classes. Operators modifying the ZZ–fermion couplings exhibit sizable interference with the SM amplitude, leading to asymmetric constraints around the SM point. In contrast, certain four-fermion operators show a more symmetric dependence, indicating a dominant contribution from the quadratic term. This behavior highlights the particular sensitivity of this channel to four-fermion operators involving bottom quarks.
The resulting 95% C.L. intervals on the WCs obtained from this procedure are summarized in Table 4. The constraints obtained in this analysis can be compared with existing global SMEFT interpretations from the CMS Collaboration [34]. In the CMS combined analysis, stringent limits are reported on operators that modify the couplings of electroweak gauge bosons to fermions, such as $C_{H\ell}^{(1,3)}$ and $C_{Hq}^{(1,3)}$, under flavor-universal or top-specific assumptions. These limits, typically at the level of $\mathcal{O}(10^{-3})$$\mathcal{O}(10^{-2})~\mathrm{TeV}^{-2}$ at 95% C.L., are primarily driven by electroweak precision observables, Higgs measurements, and top-quark processes.
@@
Figure 4: Quadratic SMEFT fits for the six operators under study. Each panel shows the fitted yield as a function of the WCs for the corresponding operator, with points representing simulated events and the red line indicating the quadratic fit. The legend shows the operator name following the SMEFTsim 3.0 conventions.
In contrast, the present analysis focuses on flavor-resolved WCs involving second-generation leptons and third-generation quarks. The constraints derived for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ can therefore be directly interpreted as channel-specific probes of the same operator classes considered in the CMS global fit, but obtained from a single decay mode and a cut-based analysis.
The constraints obtained here for $C_{H\ell}^{(1,3),22}$ and $C_{Hq}^{(1,3),33}$ are found to be of the same order of magnitude as those from the CMS global fit, despite being derived from a single decay channel and a cut-based analysis. For the four-fermion operators $C_{\ell q}^{(1,3),2233}$, which describe contact interactions between second-generation leptons and third-generation quarks, no direct constraints are currently available from global SMEFT fits, where such flavor-specific four-fermion operators are typically neglected or marginalized. The limits presented in this work therefore provide the first process-specific constraints on these operators from $Z$-boson decays. These results demonstrate the unique sensitivity of the $Z\to\mu\mu bb$ final state to four-fermion interactions involving bottom quarks, which are difficult to access in inclusive or flavor-universal analyses. The projected HL-LHC dataset, with an integrated luminosity of 3000 fb-1, further improves the constraints and providing a flavor-resolved probe complementary to existing CMS global analyses. In particular, the HL-LHC projections confirm the unique potential of the $Z\to\mu\mu bb$ channel to constrain four-fermion interactions involving bottom quarks.
@@
Table 4: 95% C.L. allowed regions for $C/\Lambda^{2}$ [TeV-2].

### v3

@@ Abstract
The Standard Model Effective Field Theory (SMEFT) provides a systematic framework to parametrize indirect effects of heavy new physics in precision measurements. In this work, we study the $\mu^{+}\mu^{-}b\bar{b}$ final state in the reconstructed $Z$-pole region and derive constraints on selected dimension-six SMEFT operators involving second-generation leptons and third-generation quarks. Signal and background processes are simulated at leading order using standard Monte Carlo tools, followed by parton showering and fast detector simulation with simplified $b$-tagging effects. We focus on four-fermion operators involving leptons and bottom quarks, as well as operators modifying effective ZZ–fermion couplings. The dependence of the selected event yield on individual Wilson coefficients is obtained using SMEFT reweighting and parametrized including both interference and quadratic dimension-six contributions. Expected constraints are derived with an Asimov likelihood-ratio approach for integrated luminosities of $138~\mathrm{fb}^{-1}$ and $3000~\mathrm{fb}^{-1}$. The results provide channel-specific constraints on flavor-resolved SMEFT interactions in mixed leptonic–hadronic $Z$-pole final states and offer a useful reference for future analyses including higher-order and systematic effects.
@@ I Introduction
The Standard Model Effective Field Theory (SMEFT) offers a model-independent framework to parameterize the effects of heavy new physics through higher-dimensional operators constructed from SM fields [4, 5, 6]. In recent years, SMEFT interpretations of $Z$-boson observables have attracted considerable attention. Purely leptonic decay modes, such as $Z\to 4\ell$, provide stringent constraints on electroweak and four-lepton operators, while complementary information has been obtained from SMEFT analyses of inclusive $Z$ production and ZZ+jets, probing quark electroweak couplings and dipole interactions [7, 8].
Despite this progress, $Z$-boson decays into mixed leptonic–hadronic final states remain relatively unexplored in the SMEFT context. Such channels are sensitive to a distinct set of operators, in particular four-fermion contact interactions involving both leptons and quarks, as well as flavor-dependent modifications of ZZ–fermion couplings [9]. Among these, the $\mu^{+}\mu^{-}b\bar{b}$ final state in the $Z$-pole region is of special interest. The presence of bottom quarks enhances sensitivity to operators coupling to the third quark generation, while the dimuon final state allows for precise reconstruction of the leptonic kinematics. At the same time, the analysis poses experimental challenges due to sizable QCD backgrounds and the reliance on heavy-flavor tagging [10, 11].
Figure 1: Representative Feynman diagram contributing to the $\mu^{+}\mu^{-}b\bar{b}$ final state near the $Z$ pole in the Standard Model, where the four-body final state may arise from final-state radiation of an off-shell $Z^{\ast}/\gamma^{\ast}$.
@@
The analysis is based on leading-order Monte Carlo simulations combined with parton showering and fast detector simulation. Constraints on the selected Wilson coefficients are obtained with an Asimov likelihood-ratio method by varying one operator at a time. Within this setup, the study provides channel-specific constraints on flavor-resolved SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state. The results also provide a useful starting point for future refinements including higher-order corrections, systematic uncertainties, electroweak interference effects, and multi-operator SMEFT interpretations.
@@ II Effective field theory description
where $\mathcal{O}_{i}$ are dimension-six operators and $C_{i}$ the corresponding WCs [8, 9]. In this work, we focus on the effects of dimension-six operators. Both the interference between the SM and dimension-six amplitudes and, in the baseline numerical results, the quadratic contribution from dimension-six amplitudes are considered. The latter is formally of order $1/\Lambda^{4}$ and should therefore be interpreted with the usual SMEFT truncation caveat.
For a generic event yield, the SMEFT dependence can be written as
$N(\vec{C})=N_{\mathrm{SM}}+\sum_{i}\frac{C_{i}}{\Lambda^{2}}N_{i}^{\mathrm{int}}+\sum_{i,j}\frac{C_{i}C_{j}}{\Lambda^{4}}N_{ij}^{\mathrm{quad}},$ (2)
where $N_{i}^{\mathrm{int}}$ denotes the interference between the SM amplitude and the amplitude induced by a single dimension-six operator, while $N_{ij}^{\mathrm{quad}}$ denotes the quadratic contribution from dimension-six amplitudes. In the numerical analysis below, one operator is varied at a time, so that Eq. (2) reduces to a quadratic dependence on a single parameter. The resulting limits are therefore individual expected constraints and should not be interpreted as fully marginalized SMEFT constraints.
@@
The selected $\mu^{+}\mu^{-}b\bar{b}$ final state near the $Z$ pole receives contributions from several classes of dimension-six operators. First, four-fermion operators involving leptons and bottom quarks generate direct contact interactions contributing to the four-body final state at tree level [7]. Schematically, such operators can be written as $(\bar{\ell}\Gamma\ell)(\bar{q}\Gamma q)$ and induce amplitudes that interfere with the SM contributions mediated by off-shell electroweak bosons.
@@
Second, operators that modify the couplings of the $Z$ boson to fermions lead to indirect contributions to the decay process. After electroweak symmetry breaking, these operators generate shifts in the effective ZZ–lepton and ZZ–quark couplings, which affect both the production and decay of the $Z$ boson [12, 13]. These effects propagate into the $Z\to\mu\mu bb$ final state through the modified electroweak vertices. Throughout the event generation and reweighting procedure, the total $Z$-boson width is kept fixed to its SM value. Operators such as $C_{H\ell}^{(1,3)}$ and $C_{Hq}^{(1,3)}$ can in general modify the total and partial $Z$ widths. The expected constraints reported in this work should therefore be interpreted under the fixed-width assumption. Including SMEFT-induced width variations would require a global treatment of electroweak precision observables and is left for future work.
@@
In the numerical analysis, we select a set of six dimension-six operators that contribute to this decay channel and study their impact individually. The operators are implemented using the SMEFTsim_general_MwScheme_UFO model [14], which is compatible with MadGraph5_aMC@NLO [15] and allows event generation and reweighting at leading order in a consistent electroweak input scheme.
In the baseline numerical results, both interference and quadratic dimension-six terms are retained in the event yield. This convention follows the structure of the squared matrix element used in the reweighting procedure and is particularly relevant for operators with suppressed interference. The associated EFT truncation caveat is discussed above and is taken into account when interpreting the quoted constraints.
The validity of the SMEFT expansion relies on the assumption that the typical energy scale of the process remains below the cutoff scale $\Lambda$ [16, 8]. Within the selected $Z$-pole region, the analysis is dominated by moderate kinematic scales set by the reconstructed invariant mass and object transverse momenta, which supports the use of the SMEFT description in the phase space considered. A simultaneous multi-operator fit is beyond the scope of the present exploratory study. Since correlations among operators can weaken individual constraints and may lead to approximate flat directions, the bounds quoted below represent individual one-operator-at-a-time expected constraints. A global treatment including correlations with other electroweak and flavor observables would be required for a complete SMEFT interpretation.
@@ III Event generation and simulation
All processes are generated at leading order using MadGraph5_aMC@NLO, followed by parton showering and hadronization with Pythia8 [17]. Detector-level effects are simulated using Delphes [18] with a CMS-like detector configuration. Unless stated otherwise, proton–proton collisions at a center-of-mass energy of $\sqrt{s}=13~\mathrm{TeV}$ are assumed throughout this analysis. All numerical results presented below should be understood as leading-order constraints within the simulation setup described here. Higher-order QCD and electroweak corrections, as well as scale and PDF uncertainties, are not included in the baseline simulation. For several of the processes considered here, next-to-leading-order predictions can be obtained with public Monte Carlo tools and may modify both the normalization and the shapes of kinematic distributions. Their inclusion is left for future precision studies.
The primary final state considered in this work is $pp\to\mu^{+}\mu^{-}b\bar{b}$ in the reconstructed $Z$-pole region. For brevity, we refer to this topology as the $Z\to\mu\mu b\bar{b}$ channel. At the amplitude level, however, the same final state may receive contributions from off-shell electroweak bosons and photon-mediated diagrams. The present analysis should therefore be understood as a fiducial sensitivity study of the $\mu^{+}\mu^{-}b\bar{b}$ final state around the $Z$ resonance rather than a fully gauge-invariant separation of individual diagram classes.
Signal events are generated at parton level using MadGraph5_aMC@NLO. Basic generator-level cuts are applied to ensure numerical stability and reflect the kinematic acceptance of the detector: $p_{T}^{\ell}>10~\mathrm{GeV}$, $|\eta^{\ell}|<2.5$ for charged leptons, and $p_{T}^{q}>10~\mathrm{GeV}$, $|\eta^{q}|<2.5$ for quarks. The resulting leading-order cross section for the selected signal topology is $\sigma=0.750~\mathrm{fb}$ at generator level before the full detector-level selection.
The reconstructed kinematic distributions shown in Fig. 2 provide an important validation of the signal modeling after detector effects are taken into account. In Fig. 2(a), the invariant mass distribution of the $b\bar{b}$ system is considerably broader than that of the dimuon system, as expected from jet energy resolution effects and QCD radiation. Nevertheless, the combined four-body invariant mass $m_{\mu\mu bb}$ displays a clear enhancement around $m_{Z}$, indicating that the signal topology can be reliably reconstructed at detector level despite the presence of hadronic final states.
@@
Figure 2: Detector-level kinematic distributions for the selected $\mu^{+}\mu^{-}b\bar{b}$ final-state topology obtained from Delphes simulation.
@@
Subleading backgrounds from electroweak processes involving photons are also considered, including $Z\gamma$, $\gamma\gamma$, and mixed $\gamma Z$ production, where the photon converts into a fermion pair. We note that electroweak amplitudes leading to the same $\mu^{+}\mu^{-}b\bar{b}$ final state can in principle interfere. In the present exploratory analysis, the samples are organized according to their dominant physical topology in order to estimate their relative impact after selection. A fully gauge-invariant treatment of the irreducible electroweak $\mu^{+}\mu^{-}b\bar{b}$ contribution, including all interferences among $Z^{\ast}$- and $\gamma^{\ast}$-mediated diagrams, is beyond the scope of this work. This point is taken into account by interpreting the numerical results as channel-specific constraints within the event-generation setup described above.
@@
Table 1: Leading-order cross sections for the signal and background processes considered in this analysis. The values are obtained at generator level before the full detector-level event selection and are intended to summarize the relative normalization of the simulated samples. The merged ZZ+jets sample, combining matrix-element calculations with up to two additional jets, is used in the final analysis.
@@ IV Event Reconstruction and Selection Strategy
Bottom-quark jets are identified using the $b$-tagging information stored in the Delphes output [18, 21]. A jet is considered a $b$-jet if its $b$-tag discriminator exceeds a threshold of $0.5$, corresponding to a typical medium working point used in CMS analyses.
@@
To suppress backgrounds containing genuine missing transverse energy, such as top-quark pair production, events are required to have $\mathrm{MET}<30~\mathrm{GeV}$. This exploits the absence of prompt neutrinos in the selected $\mu^{+}\mu^{-}b\bar{b}$ signal topology.
@@
The selection criteria are designed to preserve the characteristic kinematics of the $Z\to\mu\mu bb$ decay while suppressing backgrounds with genuine missing energy or misidentified heavy-flavor jets. For events passing all selection requirements, the invariant mass of the $\mu\mu bb$ system is reconstructed using the four-momenta of the two leading isolated muons and the two selected $b$-jets. The reconstructed invariant mass $m_{\mu\mu bb}$ serves as the primary observable for signal extraction and the subsequent SMEFT analysis [7].
The signal region is defined as $80~\mathrm{GeV}<m_{\mu\mu bb}<115~\mathrm{GeV}$, centered around the $Z$ boson mass. This window retains a large fraction of the selected signal topology while reducing non-resonant backgrounds. It provides a simple and experimentally motivated signal region for the subsequent SMEFT constraint extraction.
@@
FIG. 3(b) focuses on the signal region defined by $80<m_{\mu\mu bb}<115$ GeV. In this window, the relative contribution of the selected signal topology is enhanced, while the $t\bar{t}$ background is reduced by the missing transverse energy requirement. The remaining electroweak and associated-production backgrounds, including $Z\gamma$ and $ZH$ production, exhibit similar kinematic features but are subleading in rate after the applied selection. The invariant mass $m_{\mu\mu bb}$ is therefore used as the main observable defining the signal region for the subsequent reweighting and statistics-only constraint extraction.
V SMEFT Reweighting and Constraint Extraction
The impact of dimension-six operators on the selected $\mu^{+}\mu^{-}b\bar{b}$ final-state topology is evaluated within the SMEFT framework using an event reweighting technique [26, 27]. In this approach, the dependence of the selected event yield on each WC is obtained by reweighting Monte Carlo events. The total event weight includes the SM contribution as well as linear and quadratic corrections from dimension-six operators, allowing an efficient exploration of the WC parameter space without regenerating independent samples for every parameter point.
SMEFT effects are incorporated using the SMEFTsim_general_MwScheme_UFO model from the SMEFTsim 3.0 framework [14], which provides a gauge-invariant implementation of the relevant dimension-six operators in the $M_{W}$ input scheme. The $M_{W}$ scheme is adopted to consistently account for electroweak input-parameter shifts induced by dimension-six operators. For each generated event, a set of precomputed weights corresponding to different values of $\theta$ is stored at generator level. After parton showering, hadronization, detector simulation, and event selection, these weights are used to parametrize the SMEFT dependence of the selected yield.
@@ V SMEFT Reweighting and Constraint Extraction
where $c$ denotes the Standard Model (SM) expectation, $b$ corresponds to the interference between the SM and dimension-six amplitudes, and $a$ encodes the pure quadratic SMEFT contribution [7]. In principle, the reweighting information allows the SM contribution, the SM–dimension-six interference term, and the quadratic dimension-six contribution to be extracted separately. In the present analysis, the coefficients $a$, $b$, and $c$ are determined by fitting the reweighted event yields at several benchmark values of $\theta$. This provides a simple numerical implementation of the same quadratic dependence and is used as a consistency check of the reweighting behavior after event selection. The fitted parametrization is used only in the parameter region relevant for the quoted confidence intervals.
@@
which is compared with a $\chi^{2}$ distribution with one degree of freedom under the large-sample approximation of Wilks’ theorem [29]. The 95% confidence level (C.L.) interval on $\theta$ is obtained by requiring $q(\theta)=3.84$. No nuisance parameters are included in the baseline likelihood. The resulting intervals therefore correspond to statistics-only expected constraints within the baseline likelihood setup.
This procedure is applied independently to each dimension-six operator, with all other WCs set to zero, providing channel-specific constraints on possible SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state. In this analysis, a flavor-specific setup is adopted, in which the WCs are defined for individual fermion generations. In particular, operators involving second-generation leptons ($\ell_{2}$) and third-generation quarks ($q_{3}$) are considered explicitly. This choice provides channel-specific constraints on flavor-resolved SMEFT effects in the $\mu^{+}\mu^{-}b\bar{b}$ final state, beyond the flavor-universal assumptions commonly employed in global SMEFT fits.
@@
The observed quadratic behavior is consistent with the expected structure of the squared matrix element when one Wilson coefficient is varied at a time. It also illustrates the relative importance of linear and quadratic terms for different operator classes. Operators modifying the ZZ–fermion couplings exhibit sizable interference with the SM amplitude, leading to asymmetric constraints around the SM point. In contrast, certain four-fermion operators show a more symmetric dependence, indicating a dominant contribution from the quadratic term. This behavior highlights the particular sensitivity of this channel to four-fermion operators involving bottom quarks.
The resulting expected 95% C.L. intervals on the WCs obtained from this procedure are summarized in Table 4. These intervals are statistics-only, one-operator-at-a-time expected constraints obtained at leading order.
@@
Figure 4: Dependence of the selected signal-region yield on individual Wilson coefficients in the parameter range relevant for the expected 95% C.L. intervals. The curves are obtained from the quadratic parametrization $N(\theta)=c+b\theta+a\theta^{2}$ using SMEFT reweighted samples. Only one coefficient is varied at a time, while all other coefficients are set to zero. The results include both interference and quadratic dimension-six contributions.
A direct numerical comparison with existing global SMEFT fits should be made with care. Global analyses typically combine electroweak precision data, Higgs measurements, diboson production, top-quark observables, and differential LHC measurements. They may also adopt different flavor assumptions, operator subsets, input schemes, and SMEFT truncation prescriptions. In particular, bounds may differ significantly depending on whether they are obtained one operator at a time or after marginalization over a larger parameter space, and whether quadratic dimension-six terms are retained.
For this reason, the results in Table 4 should be viewed as channel-specific expected constraints rather than direct replacements for global SMEFT constraints. Their main value is to illustrate that the $\mu^{+}\mu^{-}b\bar{b}$ final state has potential sensitivity to flavor-resolved operators involving second-generation leptons and third-generation quarks, especially four-fermion structures that are not always explicitly resolved in flavor-universal fits. The HL-LHC luminosity scenario illustrates the statistical improvement expected at larger datasets, while a full high-luminosity interpretation would also require a dedicated treatment of systematic uncertainties.
@@
Table 4: 95% C.L. allowed regions for $C/\Lambda^{2}$ [TeV-2]. The results are obtained at leading order using the fixed-width approximation and include both interference and quadratic dimension-six contributions.
V.1 Uncertainty treatment
The likelihood used in the baseline results includes only the Poisson statistical uncertainty of the expected event yield. Experimental systematic uncertainties, such as luminosity, muon reconstruction efficiency, $b$-tagging efficiency, jet energy scale, and background normalization uncertainties, are not included. The resulting intervals therefore represent statistics-only expected constraints in the baseline setup.
A complete treatment of systematic uncertainties would require nuisance parameters constrained by auxiliary measurements. Schematically, for each bin or signal region one may write
$N_{k}(\vec{C},\vec{\eta})=N_{k,\mathrm{sig}}(\vec{C})(1+\eta_{s})+\sum_{b}N_{k,b}(1+\eta_{b}),$ (6)
with Gaussian constraints on the nuisance parameters,
$\mathcal{L}(\vec{C},\vec{\eta})=\prod_{k}\mathrm{Pois}\left(n_{k}\mid N_{k}(\vec{C},\vec{\eta})\right)\prod_{j}\exp\left(-\frac{\eta_{j}^{2}}{2\sigma_{j}^{2}}\right).$ (7)
Including such effects would generally weaken the projected bounds, especially for the HL-LHC luminosity scenario where statistical uncertainties become smaller. Since the present work is intended as a first constraint extraction, the incorporation of experimental systematics is left for future work.
In addition, dimension-six effects are included only for the selected signal topology in the baseline reweighting procedure, while the background samples are kept at their SM predictions. Operators modifying ZZ–fermion couplings may also affect electroweak backgrounds such as $ZZ$ or ZZ+heavy-flavor production. A full SMEFT interpretation should include these effects consistently together with irreducible interferences among electroweak amplitudes.